Using conversion method. Calculate the molality and molarity of aqueous phosphoric acid solution 1.88 g/ml containing 20% by mass of H3PO4

1
Expert's answer
2022-02-19T06:53:20-0500

If V = 1000 mL, we have:

m(sol) = Vd = 1000*1.88 = 1880 g

m(H3PO4) = 0.2*1880 = 376 g

n(H3PO4) = m/M = 376/98 = 3.84 mol

c(H3PO4) = 3.84/1 = 3.84 mol/L

Cm(H3PO4) = 3.84/(1.880-0.376) = 2.57 mol/kg